Minutes of the 16th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer), Peter Byrne, Deirdre Ninaber (Ladies’ Coordinator)

Absent(with apologies)

Absent

Date 4 Dec 2001 1850 – 2140

Distribution DURC Committee

Overview of topics

1 Secretarys Report 1

2 CRO Report 1

3 A&E Report 2

4 RRPC Invitational shoot 2

5 Pinks 2

6 Risk Assessment Document 3

7 DURC Clothing 3

8 AOB 3

1 Secretarys Report

8.1 Fundraiser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 8.2 Sports Campinile magazine . . . . . . . . . . . . . . . . . . . . . . . . 3 8.3 British Universities Top XX . . . . . . . . . . . . . . . . . . . . . . . . 3 8.4 World University Games . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.5 DURC 10m CAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.6 UCESSA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.7 NSRA Postals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.8 DURC 10m Shield engraving . . . . . . . . . . . . . . . . . . . . . . . 4 8.9 Club Ties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1 Secretarys Report

Mark reported the committee and ROs were formally invited to the UCDRC Christmas party on wednesday. The latest copy of the NTSA calendar was displayed, with notes on the National Squad training dates, one of which conflicted with the Colours match. Moving it to March 16/17 was discussed. It was agreed to discuss this with UCD at the christmas party. Task: Discuss change of date of colours with UCDRC committee (Steven,John)

2 CRO Report

Liam Murray has missed another RO duty. Alan wondered what we do now, in regard to disciplinary action. He also wondered if he should include Liam on next terms rota. Steven said to talk to him first, to send an email explaining the problem. Task: Talk to Liam regarding RO duties (Alan)

Alan reported on looking for the interlock for the keys. He found most switches very expensive. Mark suggested a homebrew solution, but this was thought too difficult. The model Alan had found was the type used in hotels for operating lights, and it was felt this was too easily breakable. A more industrial keyswitch is needed. Alan will source this. Mark suggested Farnell. John thought two were needed, one for .22 and one for air rifle. A discussion and field trip ensued and it was decided only one was needed, with the general illumination lights inside the range door being the point for the keyswitch. Task: Source and price hardware for keyswitch (Alan)

3 A&E Report

John reported we have 10000 more Finale Match pellets.

John also reported that Rifle V was back, but Rifle S was more badly damaged than previously thought, as it’s cylinder was out of commission. Dave Cooney has estimated £130-140 for repairs. Mark complained about the general state of the air rifles. John will inspect them.

4 RRPC Invitational shoot

Task: Inspect Air Rifles (John)

Mark pointed out that the original problem with V - the action not automatically opening during the charging cycle - was still happening, and was now happening on T as well. All were most unhappy that this had not been fixed by Dave. Steven felt that we couldn’t expect it to be repaired as we didn’t specify a general service when handing over the rifle. Noone present felt that justified not repairing the fault.

Mark repeated his complain about the general state of the rifles. A cleaning day was agreed on, and scheduled for next tuesday, December 11. John suggested bringing all the ROs as a training exercise. Mark thought some would get bored as they weren’t experienced enough to help. Steven suggested announcing it and any interested would come. Alan suggested that if cleaning didn’t help the actions to open that we send the rifles for servicing to Dave.

4 RRPC Invitational shoot

Mark announced that after some diplomacy, Steven was now invited to attend the RRPC turkey shoot. Steven then said he couldn’t make it. Mark was flummexxed.

A departure time from the range was mentioned, with John requesting it be as late as possible. No time was set however.

5 Pinks

Steven read out the proposed revised Pinks qualifications. After some discussion, they were accepted. Alan noted that the number of colours awarded each year should not depend on how many scarves we had in inventory.

6 Risk Assessment Document

John reported that DUCAC wanted this within the week. He stressed that the document could assist us in getting funding to pay for needed measures, like first aid training for ROs. It was decided to prepare this document here and this was done. Task: Type up Risk assessment document (Mark)

7 DURC Clothing

All were in favour of getting more non-shooting club clothing. John knows of a contact for T-shirts, so does Steven. Steven’s plan is that Beamish will give us T-shirts with their logo, we get them embroidered and sell them on. This gives us club T-shirts for a minimum outlay.

Mark pointed out that he wanted to get decent gear as well, and this was generally agreed on. All four present were to contact other clubs and find out their arrangements and costs.

8 AOB

Task: Find out about other club’s arrangements for club clothing (Mark,John,Alan,Steven)

8 AOB

8.1 Fundraiser

Alan wanted to run a Pub Quiz as a club fundraiser. John suggested using the Pav and getting Guinness sponsorship. Steven recommended talking to Liam as he was running one at present. Mark wondered if Guinness would give Tshirts, Alan wondered if they’d give beer. John explained the convoluted approach to guinness sponsorship in college at present. It was agreed John would attemt to secure beer for this event. Task: Arrange guinness beer for quiz (John) Task: General logistics for quiz (Alan)

8.2 Sports Campinile magazine

This new newsletter from DUCAC was raised by Steven. Mark and John had heard about it, but too late to do anything for it. John said only high-profile stuff should go into it. Mark had a strong objection to the idea of not trying continously to get articles into any kind of press. Steven agreed to hold a review before each issue where possible, to select items for publication, if any were available. Task: Hold review in feburary for articles (Steven)

8.3 British Universities Top XX

Steven had been on the phone to BUSA and the NSRA. The Top XX is now run by BUSA with technical support only from the NSRA, and BUSA were only sending entry forms to affiliated universities. As we’re not affiliated we won’t be able to enter. John felt it wasn’t a big deal to get TCD to affiliate, and agreed to bring it up at the next DUCAC meeting. Task: Discuss BUSA affiliation at next DUCAC meeting (John)

8.4 World University Games

Steven also reported that the NSRA and BUSA were disagreeing over this - BUSA didn’t want to send a UK team, the NSRA did. This will be resolved in a few weeks. Steven though, is ineligible to represent Ireland.

8.5 DURC 10m CAR

Mark handed round the CAR for the DURC 10m Air Open.

8 AOB

8.6 UCESSA

Mark reported Dave Gallagher had missed his second round and was now disqualified and owed us his entry fee. Mark felt this should be charged seperately for disciplinary reasons (as this was the original reasoning). John felt it should just be put on his RO credits, but was overruled.

Mark reminded everyone that the next round was due on Dec 17.

Mark reported that Alan and Shane Kelly had not submitted cards for the Air Rifle postal. Alan protested, as he had shot his cards at the DURC Open. Mark said he had told Alan that the cards were lost and as such had to be reshot, that thre people (Alan, Mark and Liam Murray) had been caught by this and two had reshot.

Mark reminded everyone that the next round was due early next year and they had christmas off.

8.7 NSRA Postals

Mark went through the NSRA postal competitions and took some details for people shooting. No teams were felt to be viable due to the low number of shooters except in 10m air rifle. Mark and Steven will shoot on a team with one other (yet to be selected). Mark and Steven will also shoot 3P 25yard, and will order their own targets to do so. John and alan will also do the 25yard prone postal. Mark will continue to recruit.

8.8 DURC 10m Shield engraving

Steven agreed to get the Simon Kennedy Shield engraved for Nicola. Task: Engraving on Air Rifle Shield (Steven)

8.9 Club Ties

The club ties idea was discussed, but it was reluctantly agreed that we can’t afford them just yet (£300 is too much when we can’t yet afford to buy ammo in bulk).

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

8 AOB

Tasks agreed on at meeting

Steven,John Discuss change of date of colours with UCDRC committee (4 Dec

2001) [/138] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Alan Talk to Liam regarding RO duties (4 Dec 2001) [/147] . . . . . . . . /2 Alan Source and price hardware for keyswitch (4 Dec 2001) [/161] …. /2 John Inspect Air Rifles (4 Dec 2001) [/173] . . . . . . . . . . . . . . . . . /2 Mark Type up Risk assessment document (4 Dec 2001) [/217] . . . . . . . /2 Mark,John,Alan,Steven Find out about other club’s arrangements for club clothing (4 Dec

2001) [/233] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

John Arrange guinness beer for quiz (4 Dec 2001) [/247] . . . . . . . . . . /2 Alan General logistics for quiz (4 Dec 2001) [/248] . . . . . . . . . . . . . /2 Steven Hold review in feburary for articles (4 Dec 2001) [/260] . . . . . . . /2 John Discuss BUSA affiliation at next DUCAC meeting (4 Dec 2001) [/272] /2 Steven Engraving on Air Rifle Shield (4 Dec 2001) [/321] . . . . . . . . . . /2